OpenStudy (anonymous):

find two numbers whose difference is 22 and whose product is as small as possible

zepdrix (zepdrix):

Hey Ivan :) Welcome to OpenStudy! What type of class is this for? Can we use calculus techniques?

OpenStudy (anonymous):

college algebra, and yes any technique works...

zepdrix (zepdrix):

Ooo boy :( I'm trying to remember how to do optimization problems without using Calculus. hmmm

zepdrix (zepdrix):

Oh oh oh ok ya ya ya, we can do this.

zepdrix (zepdrix):

The difference of two numbers is 22.\[\Large\rm x-y=22\]Their product is a minimum.\[\Large\rm P_{\min}=xy\]I'm calling it P for product. If we use our first equation, add y to each side,

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{x=22+y}\]

zepdrix (zepdrix):

We can plug this into our product equation:\[\Large\rm P_{\min}=\color{orangered}{x}y\]\[\Large\rm P_{\min}=\color{orangered}{(22+y)}y\]

zepdrix (zepdrix):

Distribute,\[\Large\rm P_{\min}=y^2+22y\]So what we have here is a parabola. You have to remember a little something about parabolas in order to solve this equation. The leading term is positive, so it's opening "up". So our minimum is going to be the `vertex` of this parabola.

zepdrix (zepdrix):

So we need it in vertex form.

OpenStudy (anonymous):

difference is 22 and product is as mall as possible?